OpenStudy (anonymous):

How do you solve this linear equation using substitution? The decimals keep messing me up. [0.5x+3.2y=9.0] [0.2x-1.6y=3.6]

OpenStudy (anonymous):

Start with .5x + 3.2y = 9, and solve for x. .5x = 9 - 3.2y x = (9 - 3.2y)/2 = 4.5 - 1.6y Then substitute this value in for x in your second equation. .2 ( 4.5 - 1.6y) = 3.6 .9 - .32y = 3.6 Solve for y. Then put whatever number you get for y into your x = 4.5 - 1.6y, to find your x value.

OpenStudy (asnaseer):

if you are not comfortable with decimals, you can always multiply both sides of the equation by a power of ten to get rid of them. e.g. 3.2 * 10 = 32 3.23 * 100 = 323 so in your equation you have:\[0.5x+3.2y=9.0\]\[0.2x-1.6y=3.6\]if you multiply each of these equations by 10 they become:\[5x+32y=90\]\[2x-16y=36\]no more decimals! I hope this helps you...